OpenStudy (anonymous):

Can someone please help me with this? I had to solve it once before but I have to turn in the notes to my teacher and I can't find them and right now I can't remember how to work the formula, so I'm having some issues. Help would be greatly appreciated! Thank you! :) Solve each of the quadratic equations below and describe what the solution(s) represent to the graph of each. Show your work to receive full credit. 0 = x2 + 5x + 6 0 = x2 + 4x + 4

OpenStudy (blurbendy):

Ill do the first, and hopefully you can do the second: x^2 +5x +6 = 0 (x+2)(x+3) = 0 x + 2 = 0 x = -2 x + 3 = 0 x = -3

OpenStudy (anonymous):

http://www.analyzemath.com/Equations/Quadratic-1.html for the second equation the D=0 this means that there will be only one zero

OpenStudy (anonymous):

and that zero will be -b/(2a) in this case -(1/2)

OpenStudy (anonymous):

Thank you blurbendy! :) Thank you Noided! :) Also, Noided, I'm really confused... I'm not very good at quadratics, so I'm not exactly sure how to do step two. I need to group them, right?

OpenStudy (anonymous):

hmm wait

OpenStudy (anonymous):

step1 is calculating the discriminant, right?

OpenStudy (anonymous):

I think so, yes. lol

OpenStudy (anonymous):

then step 2 would be deciding what to do D<0 don't do anything D=0 calculate -b/(2a) if D>0 calculat (-b+sqrt(D))/(2a) and also (b+sqrt(D))/(2a)

OpenStudy (anonymous):

okidokie pianissimokie?

OpenStudy (anonymous):

easy peasy?

OpenStudy (anonymous):

you get it Maddy? It's easy now huh?!

OpenStudy (anonymous):

this can give you some insights, play around a bit with it: http://www.mathsisfun.com/quadratic-equation-solver.html if D=0 the top of the function will touch the x-axis D>0 it will cross it twice D<0 it will never cross nor touch it

OpenStudy (anonymous):

Uhm, honestly I'm still really confused. lol :/ I suck at math.

OpenStudy (blurbendy):

personally, it's easier for me to factor first, then just solve for x. that way you don't have to worry about solving the discriminant. did you follow my first example?

OpenStudy (anonymous):

I was taught to factor first as well, that might be why I'm getting confused. I'm trying to follow your first example.

OpenStudy (blurbendy):

basically i found which two numbers ADD up to 5 and MULTIPLY to get 6

OpenStudy (anonymous):

Is this correct? 0=x^2+4x+4 x^2+4x+4=0 (x+2)^2+0 x+2=0 x=-2

OpenStudy (blurbendy):

perfect

OpenStudy (anonymous):

Okay, thank you guys soo much!!! I really appreciate it!!! :)

OpenStudy (blurbendy):

anytime!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

I did the formula's with a friend last time and I couldn't remember how to do it on my own. If I could've done the problems on my own, I would have but I needed some help. I don't understand quadratic formula's and I kinda suck at math. So, they weren't excuses, it was factual. Have a good day. :)